OpenStudy (anonymous):

How many solutions does the following system of equations have? y+5/2 x+2 2y=5x+4 A. Two B. One C. Zero D. Infinitely many

OpenStudy (anonymous):

y+5/2 x+2 2y=5x+4 y=-5/2x+2 y=5/2x+2 since the slopes are different, one solution, B

OpenStudy (anonymous):

it was D :/

OpenStudy (anonymous):

look at your first equation; i misread it but there is no equals sign; if it is y=5/2x+2 then i agree with answer D. sorry I misread it; maybe getting a little tired to read a + as an =